OpenStudy (anonymous):

Algebra 2 : c( x + 2) - 5 = b( x - 3)

HanAkoSolo (jamierox4ev3r):

just distribute it, u should get an answer c(x+2)-5=b(x-3) cx+2c-5=bx-3b

HanAkoSolo (jamierox4ev3r):

from there u should be able to figure it out @Stenly